OpenStudy (anonymous):

Of 5 employees, 3 are to be assigned an office and 2 are to be assigned a cubicle. If 3 of the employees are men and 2 are women, and if those assigned an office are to be chosen at random, what is the probability that the offices will be assigned to 2 of the men and 1 of the women?

OpenStudy (anonymous):

\[required probability=3C2*2C1=\frac{ 3*2 }{ 2*1 }*\frac{ 2 }{ 1 }=6 ways \]